OpenStudy (anonymous):

An experiment was conducted in which the initial concentration of C2H5Cl was 8.80 x 10-5 M. At the temperature of the experiment, the reaction had a rate constant of 7.41 x 10-3 s-1. How long would it take for the concentration of C2H5Cl to be reduced to one eighth (1/8) of its original concentration?

OpenStudy (aaronq):

What order is the reaction?

OpenStudy (anonymous):

@aaronq the reaction is first order sorry I forgot to include it

OpenStudy (aaronq):

This is just exponential decay, you can do it in various ways. The easiest is using the integrated rate-law equation: \(ln[A]_t=-kt+ln[A]_o\)

OpenStudy (aaronq):

so set \([A]_o=8.80 *10^{-5} M\) and \([A]_t=\dfrac{8.80 *10^{-5} M}{8}\)

OpenStudy (anonymous):

So, the reaction is a first order one due to the rate constant ("s^-1"). The integrated equation would be \[[A]=[A ]_{0}e ^{-kt}\] that once we inssert the values:\[[8.8x10^{-5}]=[1.1x10^{-5}]e ^{-(7.41x10^{-3})t}\]

OpenStudy (anonymous):

we get t=280.62 s

OpenStudy (aaronq):

it works both ways

OpenStudy (aaronq):

but the set up above is incorrect,it should be \(\large \dfrac{8.8*10^{-5}}{8}=8.8*10^{-5}*e^{-kt}\)

OpenStudy (anonymous):

Yes, sorry, introduced it wrong, but the answer is the same it should be \[[1.1x10^{-5}]=[8.8x10^{-5}]e ^{-(7.41x10^{-3})t}\] which gives us t=280.62 s